OpenStudy (anonymous):

y-y1=m(x-x1) solve for m

OpenStudy (anonymous):

divide both sides by \(x-x_1\)

OpenStudy (anonymous):

could you show me step by step?

OpenStudy (anonymous):

it is only one step

OpenStudy (anonymous):

\[y-y_1=m(x-x_1)\iff \frac{y-y_1}{x-x_1}=m\]

OpenStudy (anonymous):

oh i see now thanks